please help!
Given 2 similar tirangels, Triangle ABC and Triangle DEF, if AB=7 and DE=12 and AC=8, then DF = _____. (round to the nearest tenth.)
A. 4.7
B. 13.7
C. 14.0
Thank you!

Answer by tutorcecilia(2152)

You are working with similiar triangles which is the same as working with proportions. Compare like-sides:
7/8 = 12/x
7x=(8)(12) [Cross-multiply]
7x=96
x=13.71